an electron and a proton enter a magnetic field with same velocity then their acceleration will be?

Grade:Select Grade

1 Answers

KAPIL MANDAL
9 years ago
 
The force acting on them is F= e(v×B). Now, charges of electron and proton have opposite sign. Again from Newtons law, F=ma where a is the acceleration. So, 
   e(v×B)=ma. 
 a=e(v×B)/m
Now, charges of electron and proton have opposite sign, So,
firstly, the direction of them will be opposite.
Secondly, mass m of proton is 1836 times greater than that of the electron. So, a=magnitude of acceleration of proton will be 1836 times smaller. Ok?